Mike wants to save $47000 through semi annual payments over the next 6 years. If the interest rate is j2 = 7.1%, complete the first three lines of the sinking fund schedule. NOTE: Round the deposit appropriately and use this value in further calculations. Schedule:

Schedule:

Deposit number Interest Deposit Increase Amount
0
1
2
Answer #1

Answer:

Sinking fund is a fund created out of fixed payments each period to accumulated to a future sum after a specified period.

In our case payment is done semi annually and also interest is provided semi annually.

So formula for amount to be paid regularly is :

A = Fn * (i/[(1+i)n-1])

where Fn=$47000

i = 0.071

n= 6*2 = 12

A = 47000*(0.071/[(1+0.071)12 - 1]) = $2612 to be paid semi annually.

Lets provide the details of cash flow:

Deposite number Year Principal (Increase Amount) Semi annual Interest @ 7.1% (Interest Deposite)
0 0.5 2612.00 185.45
1 1.00 5409.45 384.07
2 1.50 8405.52 596.79
3 2.00 11614.32 824.62
4 2.50 15050.93 1068.62
5 3.00 18731.55 1329.94
6 3.50 22673.49 1609.82
7 4.00 26895.31 1909.57
8 4.50 31416.87 2230.60
9 5.00 36259.47 2574.42
10 5.50 41445.89 2942.66
11 6.00 44388.55 3151.59
Total 47540.14